OpenStudy (anonymous):

The volume of a cylinder is 81 cubic feet. What is the volume of a cone that fits exactly inside the cylinder? Provide an explanation and proof for your answer to receive full credit. @luckycoins888 @Luigi0210 @EclipsedStar @pooja195 @carsonsi HELPP PLEASE!!!

OpenStudy (campbell_st):

here are the volume formulae see if you can work out what to do. Both the cone and cylinder have the same radius and height Cylinder \[V = \pi r^2h\] Cone \[V = \frac{1}{3} \pi r^2 h\] hope it helps

OpenStudy (anonymous):

is that the anwer?

OpenStudy (campbell_st):

no. look at the 2 equations, what needs to happen to the cylinder's volume to get the cone's volume...

OpenStudy (anonymous):

tbh idk im lost I need help on this question

OpenStudy (anonymous):

@AnswerMyQuestions @Conqueror

OpenStudy (anonymous):

@Gabylovesyou

OpenStudy (anonymous):

idk :( @campbell_st

OpenStudy (campbell_st):

it sis simply a case the volume of the cone is 1/3 of the cylinder.

OpenStudy (anonymous):

im so lost!!! :( @campbell_st

OpenStudy (campbell_st):

ok... volume of the cylinder is \[V_{cylinder} = \pi r^2 h\] Volume of the cone is \[V_{cone} = \frac{1}{3} \pi r^2 h\] both equations contain \[\pi r^2 h\] and the radius and height are identical then \[V_{cone} = \frac{1}{3} \times V_{cylinder}\] so now substitute the volume of the cylinder you have been given and calculate the volume of the cone.

OpenStudy (anonymous):

yea ik that @campbell_st

OpenStudy (anonymous):

sorry i missed this

OpenStudy (anonymous):

its ok @carsonsi